# Forecasting Company Fundamentals

### v1

@@ 4.2 Realisitc Market Evaluation
We built simulated portfolios from the CF forecasts based on several different strategies. First of all, we considered two different proxies to identify good stocks, namely, the highest Operating Income or highest Total Revenues four quarters in the future, each normalized by the current Enterprise Value of the company. Secondly, we rebalanced our portfolio either after each quarter or at the start of each new year. In addition to the criteria in Section 4.1.2, we excluded the real estate and financial sectors due to special reporting semantics.To ensure our method is not merely exploiting some current sector-specific performances, we ensure our allocation replicates the sector composition of the MSCI World reference index. We similarly approximate its regional distribution as closely as possible. Our portfolios each consist of 50 stocks with a fixed 2% of the capital each. At each rebalancing step, the best possible companies based on the above criteria are chosen, and stocks are sold and bought to reflect that reevaluation. We compare our method to the reference index and two other investment strategies: First, we replicate the Clairvoyant portfolios based on the same selection criteria but using the usually unknowable exact ground truth future (Alberg & Lipton, 2018). Secondly, we compare portfolios using our forecasts with ones based on available human analysts’ expectations for total revenues.
@@
The performance of portfolios based on CF forecasts and their baselines are shown in Table 3. We can firstly see that for Operating Income, rebalancing each year is superior to each quarter, and vice versa for Total Revenues, possibly due to the special annual cycle of financial reporting also visible in Figure 6. Secondly, we can confirm that the hypothetical Clairvoyant models are effective tools for portfolio optimization, further motivating the pursuit of improving CF forecasting. Thirdly, human analysts’ predictions are highly effective for investing, possibly reinforced through self-fulfilling prophecies and feedback loops (Sant & Zaman, 1996; Mauboussin, 2002). This confirms the empirical findings of Guerard et al. (2015). Lastly, portfolios based on automated GRU forecasts are useful for investing, enabling a final portfolio value that is 10 percentage points larger than the reference index after about ten years when using the Operating Income and rebalancing yearly. While the volatility and Beta are expectedly higher than the reference index, they are on par with the other strategies. Detailed inspection of the entire time span reveals that our method consistently finds better stocks than when following the reference index in the described setting, except from the onset of the pandemic effects in Q2 2020 on to the end of 2020. Similarly, the increased volatility originates to a large degree from the time of 2019 onwards, and was comparably low for the previous years. This, again, matches with the findings in Figure 5, where pandemic uncertainty leaks back into the forecasts starting 2019 Q2. This strongly motivates going beyond the probabilistic forecasting performed so far and investigating methods to maintain reasonable performance in such unstable times too.
@@ 5 Interpretability, Explanations and Domain Expertise
When building models, it is most often worthwhile to leverage the expertise of practitioners regularly working with the data at hand (Collopy et al., 2001; Kelly & Xiu, 2023, Sec. 1.5). We thus provide an overview of which of the investigated models can reasonably benefit from such information. Most models support simple inductive biases by fixing, bounding, or regularizing certain parameters. However, this is only useful in practice where such weights can be identified with a specific piece of domain knowledge and, therefore, not helpful in most deep architectures. Therefore, it might be necessary to fall back to more general methods of incorporating human insights like explainable interactive learning (XIL), where experts improve the model by providing feedback based on explanations of model outputs (Teso & Kersting, 2019; Kraus et al., 2024). We will discuss the possibilities of applying interpretability methods to the models at hand to permit such approaches in the first place. This furthermore allows some degree of validation of the modeling outcomes and is therefore highly desirable in most practical ML applications (Collopy et al., 2001; Hoang & Wiegratz, 2023). Also, parameters such as the lookback window or forecast horizon are always opportunities to use expert knowledge and were already discussed in Section 4.1.3. In this section, we thus focus on more high-level and human-understandable aspects specific to each model.
@@
While Linear Regression permits constraining specific weights, this is hardly desirable as this model is seldom used as more than a baseline. For most other global models, opportunities for expert involvement are limited. However, general black-box explainability methods like SHAP or Integrated Gradients apply to all deep learning methods (Marcinkevičs & Vogt, 2023), although the exact insight gained from them can be challenging to understand for domain experts. Some of the models provide specific white-box methods to gain insights into the forecasts. For example, models including and building on Transformers (like TFT) allow inspecting the attention map to learn what the model “looks at” (Vaswani et al., 2017) and when attention maps are atypical (Lim et al., 2021). Furthermore, TFT allows inspecting the importance of the input variables. This is shown exemplarily in Figure 7 for all covariates with at least 3%, and static variables with at least 10% importance, respectively. The stacked N-BEATS architecture allows for modeling different task-specific basis functions in different stacks and inspecting the backcasting results. Thus, one can define detrending and seasonality-removing blocks and then inspect and validate their behavior. The same is possible for the hierarchical extension N-HiTS. In summary, all of these methods allow for interpreting the decisions and explaining the output of deep learning models. To eventually improve them, human-in-the-loop methods like XIL can be used, helping to build better and more reliable forcasters.
@@ 6 Conclusion
We established the need for an in-depth comparison of company fundamentals forecasting models in both qualitative and quantitative terms. This is relevant not only for finance but also as a case study for machine learning and econometrics. We started with an in-depth analysis of the data and solutions to its specific challenges. We explored the applicability of 22 classical statistical and modern machine learning models and identified their properties and capabilities. Subsequently, we have shown that identifying the best overall model requires a nuanced analysis. Particularly, we demonstrated that surpassing simple baselines in deterministic forecasts is challenging. Yet, the much more helpful probabilistic forecasts can be obtained by either AutoARIMA or, equivalently, by any of the deep learning models. These global models can learn from the time series of many companies, effectively estimating a more adequate distribution. We have shown that enriching the data with static context provides negligible benefits, suggesting that the key predictive patterns are already latent in the covariate and target variables. Depending on the field of application and the variables of interest, vastly different models should be selected. For example, a varying set of models excelled before, during, and after the COVID-19 pandemic, respectively. We further find the quality of the obtained forecasts to be similar to human analyst predictions. We continue by evaluating the best automated forecasts on factor models to assess their practical use in investing. We find that they can outperform the reference index over ten years and observe the benefit to be greatest in times of higher financial stability. We finally motivated and showed possibilities for interpreting the models’ behavior and generating explanations for their forecasts. This, in turn, permits domain experts to elevate and validate the quality of predictions and eventually improve them.
@@
The greatest challenge to CF forecasting seems to be the large variability of patterns across the years. This is amplified by the relatively small amount of data available and the inherent out-of-sample task. Therefore, ensemble methods and incorporating additional signals and domain knowledge appear to be very fruitful. For instance, one could mine textual features from news, investor calls, or social media by extracting sentiment scores or semantic information from that. Furthermore, causal models might help distinguish spurious correlations, possibly from data collection artifacts, from actual causal relations, like the onset of disruptive global events. This is closely tied to mining patterns in CF data for insights into the macroeconomic behavior of entire sectors or countries in relation to each other. The task is, furthermore, inherently continual since a new set of fundamentals is released each quarter. Explicit lifelong learning methods might aid in regulating which of the learned patterns still apply in the new time period.
@@ Appendix A Appendix
This supplement provides a discussion of wider ramifications of this work (Section A.1), detailed statistics on the CF dataset, highlighting the challenges of the forecasting task (Section A.2), details for reproducing the model training (Section A.3), insights on the effect of RevIn (Section A.4), and the complete results for the comparison of human analyst expectations to the model’s forecasts (Section A.5).
@@ A.2 Charactersitics of the Dataset
Even after the normalization described in Section 4.1.2, the data exhibits significant skewness and heavy tails, effectively making it considerably non-Gaussian. That is shown in Figure 8 and Table 4. This reoccurring problem in finance (Schmitt et al., 2013) invalidates some model assumptions, like of the $\operatorname{ARIMA}$-family models. This might explain some of the observed challenges with forecasting the CFs, especially using those models.
Figure 8: The histogram of the five target features. It is apparent that the data is not Gaussian.
@@
Figure 9: Historgram over the strongest seasonalities in each feature. Notably, very different seasonalities are present in the data, even for the same feature. Moreover, there is no strong seasonality after four quarters (one year), yet for all other multiples of four.
@@ A.3 Details on Model Configurations and Training
We based our analysis on the implementation in the darts software library (Herzen et al., 2022). We used scikit-learn (Pedregosa et al., 2011) for data normalization. For evaluation, we used torchmetrics (Nicki Skafte Detlefsen et al., 2022) and the CRPS implementation of Pyro (Bingham et al., 2019). Some models obtain probabilistic forecasts via specialized architectures or sampling methods, while others simply learn to predict the parameters of a distribution instead of the data directly. This is the case for Linear Regression as well as all deep learning models, where we, therefore, performed quantile regression on $0.01,0.05,0.1,0.15,0.2,0.25,0.3,0.4,0.5,0.6,0.7,0.75,0.8,0.85,0.9,0.95,0.99$.
@@
For the linear regression and random forest models, we estimated separate models for each of the target variables and four future time steps. For the latter, each forest contained 100 trees learned from bootstrapped subsamples.
We trained all deep learning models with gradient descent on a batch size of 64 for 100 epochs. We adjusted for different training duration requirements by early stopping after no change of the validation loss after three epochs. We employed the AdamW optimizer (Loshchilov & Hutter, 2018) with learning rate $10^{-4}$, weight decay weighting of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. Furthermore, we clipped gradients to $1.0$ to stabilize training. For all models except DLinear and NLinear, we used a 10% dropout rate. The following lists the specific hyperparameters for each of the deep learning architectures. For DLinear, we used a kernel size of 10 to estimate the moving average of the trend. All recurrent neural networks (GRU, LSTM, and block variants) were three layers deep. The single models and block variants had a hidden size of 64 and 128, respectively. The TCN convolutions were set to a kernel width of 3 steps, 16 filters, and a dilation of 2 time steps. The Transformer was trained with a token size of 120, a feedforward dimension of 512, GELU activations, four encoder layers, four decoder layers, and six attention heads. TFT used a hidden size of 36 and six attention heads covering the entire time span. We learned a single block of six layers with hidden dimension 512 for N-BEATS. The N-HiTS models consisted of three stacks, each with a single two-layer block and 512 hidden dimensions. Our configuration of TiDE had two encoder and decoder layers each, a hidden size of 128, and a decoder with a hidden size of 32 and an output dimension of 16.
A.4 The Benefit of Reversible Instance Norm
@@
Table 5: A comparison showing the benefit of RevIN in percentage over the models without RevIN. The scores are the nCRPS obtained from probabilistic forecasts, where lower is better.
@@
A.5 Algorithmic and Human Forecasts
To facilitate a relation of the human evaluation to the automated forecast, we provide comparable results in Table 6, as discussed in Section 4.1.3. Here, we made sure to test the models only on the companies where expert forecasts were available. Furthermore, only the fourth (one-year-ahead) time step is considered, which aligns with the human expectations. For simplicity, we show the results for an entire company in the model evaluation, even if individual human analyst predictions are missing since this happens only rarely. Since the expectations contained some heavy outliers, we omitted metric results larger than four standard deviations from the statistics. The comparison is performed on the transformed data as laid out in Section 4.1.2. Consistent with the prior evaluation, the provided nCRPS score for the deterministic analyst expectations is the special case of the MAE.

### v2

@@ 4.2 Realisitc Market Evaluation
We built simulated portfolios from the CF forecasts based on several different strategies. First of all, we considered two different proxies to identify good stocks, namely, the highest Operating Income or the highest Total Revenues for the four quarters in the future, each normalized by the current Enterprise Value of the company. Secondly, we rebalanced our portfolio either after each quarter or at the start of each new year. In addition to the criteria in Section 4.1.2, we excluded the real estate and financial sectors due to special reporting semantics. To ensure our method is not merely exploiting some current sector-specific performances, we ensure our allocation replicates the sector composition of the MSCI World reference index. We similarly approximate its regional distribution as closely as possible. Our portfolios each consist of 50 stocks with a fixed 2% of the capital each. At each rebalancing step, the best possible companies based on the above criteria are chosen, and stocks are sold and bought to reflect that reevaluation. We compare our method to the reference index and two other investment strategies: First, we replicate the Clairvoyant portfolios based on the same selection criteria but using the usually unknowable exact ground truth future (Alberg & Lipton, 2018). Secondly, we compare portfolios using our forecasts with ones based on available human analysts’ expectations for total revenues.
@@
The performance of portfolios based on CF forecasts and their baselines is shown in Table 3. We can firstly see that for Operating Income, rebalancing each year is superior to each quarter, and vice versa for Total Revenues, possibly due to the special annual cycle of financial reporting also visible in Figure 6. Secondly, we can confirm that the hypothetical Clairvoyant models are effective tools for portfolio optimization, further motivating the pursuit of improving CF forecasting. Thirdly, human analysts’ predictions are highly effective for investing, possibly reinforced through self-fulfilling prophecies and feedback loops (Sant & Zaman, 1996; Mauboussin, 2002). This confirms the empirical findings of Guerard et al. (2015). Lastly, portfolios based on automated GRU forecasts are useful for investing, enabling a final portfolio value that is 10 percentage points larger than the reference index after about ten years when using the Operating Income and rebalancing yearly. While the volatility and Beta are expectedly higher than the reference index, they are on par with the other strategies. Detailed inspection of the entire time span reveals that our method consistently finds better stocks than when following the reference index in the described setting, except for the onset of the pandemic effects in Q2 2020, through to the end of 2020. Similarly, the increased volatility originates to a large degree from the time of 2019 onwards, and was comparably low for the previous years. This, again, matches with the findings in Figure 5, where pandemic uncertainty leaks back into the forecasts starting 2019 Q2. This strongly motivates going beyond the probabilistic forecasting performed so far and investigating methods to maintain reasonable performance in such unstable times, too.
@@ 5 Interpretability, Explanations and Domain Expertise
When building models, it is most often worthwhile to leverage the expertise of practitioners regularly working with the data at hand (Collopy et al., 2001; Kelly & Xiu, 2023, Sec. 1.5). We thus provide an overview of which of the investigated models can reasonably benefit from such information. Most models support simple inductive biases by fixing, bounding, or regularizing certain parameters. However, this is only useful in practice where such weights can be identified with a specific piece of domain knowledge and, therefore, not helpful in most deep architectures. Therefore, it might be necessary to fall back to more general methods of incorporating human insights like explainable interactive learning (XIL), where experts improve the model by providing feedback based on explanations of model outputs (Teso & Kersting, 2019; Kraus et al., 2024c). We will discuss the possibilities of applying interpretability methods to the models at hand to permit such approaches in the first place. This furthermore allows some degree of validation of the modeling outcomes and is therefore highly desirable in most practical ML applications (Collopy et al., 2001; Hoang & Wiegratz, 2023; Steinmann et al., 2024). Also, parameters such as the lookback window or forecast horizon are always opportunities to use expert knowledge and were already discussed in Section 4.1.3. In this section, we thus focus on higher-level and human-understandable aspects specific to each model.
@@
While Linear Regression permits constraining specific weights, this is hardly desirable, as this model is seldom used as more than a baseline. For most other global models, opportunities for expert involvement are limited. However, general black-box explainability methods like SHAP or Integrated Gradients apply to all deep learning methods (Marcinkevičs & Vogt, 2023), although the exact insight gained from them can be challenging to understand for domain experts. Some of the models provide specific white-box methods to gain insights into the forecasts. For example, models including and building on Transformers (like TFT) allow inspecting the attention map to learn what the model “looks at” (Vaswani et al., 2017) and when attention maps are atypical (Lim et al., 2021). Furthermore, TFT allows inspecting the importance of the input variables. This is shown exemplarily in Figure 7 for all covariates with at least 3%, and static variables with at least 10% importance, respectively. The stacked N-BEATS architecture allows for modeling different task-specific basis functions in different stacks and inspecting the backcasting results. Thus, one can define detrending and seasonality-removing blocks and then inspect and validate their behavior. The same is possible for the hierarchical extension N-HiTS. In summary, all of these methods allow for interpreting the decisions and explaining the output of deep learning models. To eventually improve them, human-in-the-loop methods like XIL can be used, helping to build better and more reliable forecasters.
@@ 6 Conclusion
We established the need for an in-depth comparison of company fundamentals forecasting models in both qualitative and quantitative terms. This is relevant not only for finance but also as a case study for machine learning and econometrics. We started with an in-depth analysis of the data and solutions to its specific challenges. We explored the applicability of 24 classical statistical and modern machine learning models and identified their properties and capabilities. Subsequently, we have shown that identifying the best overall model requires a nuanced analysis. Particularly, we demonstrated that surpassing simple baselines in deterministic forecasts is challenging. Yet, the much more helpful probabilistic forecasts can be obtained by either AutoARIMA or, equivalently, by any of the deep learning models. These global models can learn from the time series of many companies, effectively estimating a more adequate distribution. We have shown that enriching the data with static context provides negligible benefits, suggesting that the key predictive patterns are already latent in the covariate and target variables. Depending on the field of application and the variables of interest, vastly different models should be selected. For example, a varying set of models excelled before, during, and after the COVID-19 pandemic, respectively. We further find the quality of the obtained forecasts to be similar to human analyst predictions. We continue by evaluating the best automated forecasts on factor models to assess their practical use in investing. We find that they can outperform the reference index over ten years and observe the benefit to be greatest in times of higher financial stability. We finally motivated and showed possibilities for interpreting the models’ behavior and generating explanations for their forecasts. This, in turn, permits domain experts to elevate and validate the quality of predictions and eventually improve them.
@@
The greatest challenge to CF forecasting seems to be the large variability of patterns across the years. This is amplified by the relatively small amount of data available. Data upsampling techniques might permit more fine-grained forecasts once higher-frequency data of at least some companies becomes available, from which interpolations might be learnable (Ghosh et al., 2021). The data scarcity and the inherent out-of-sample task makes ensemble methods and incorporating additional signals and domain knowledge appear to be very fruitful. For instance, one could mine textual features from news, investor calls, or social media by extracting sentiment scores or semantic information. Furthermore, causal models might help distinguish spurious correlations, possibly from data collection artifacts, from actual causal relations, like the onset of disruptive global events. This is closely tied to mining patterns in CF data for insights into the macroeconomic behavior of entire sectors or countries in relation to each other. The other way around—zooming in on specific companies—might reveal particularities about individual companies or their relation to other market participants. Company fundamentals forecasting is, furthermore, inherently continual since a new set of fundamentals is released each quarter. Explicit lifelong learning methods might aid in regulating which of the learned patterns still apply in the new time period.
@@
Additionally, we thank the anonymous reviewers whose comments helped improve this manuscript.
@@ Appendix A Appendix
This supplement provides a discussion of wider ramifications of this work (Section A.1), detailed statistics on the CF dataset, highlighting the challenges of the forecasting task (Section A.2), details for reproducing the model training (Section A.3), insights on the effect of RevIn (Section A.4) and providing static context (Section A.5), as well as the comparison of human analyst’s to the model’s forecasts (Section A.6).
@@ A.2 Charactersitics of the Dataset
Even after the normalization described in Section 4.1.2, the data exhibits significant skewness and heavy tails, effectively making it considerably non-Gaussian. That is shown in Figure 8 and Table 4. This reoccurring problem in finance (Schmitt et al., 2013) invalidates some model assumptions, like of the $\operatorname{ARIMA}$-family models. This might explain some of the observed challenges with forecasting the CFs with those models.
Figure 8: The five target features are substantially non-Gaussian even after normalization.
@@
Figure 9: Very different seasonalities are present in the data, even for the same feature. Moreover, there is no strong seasonality after four quarters (one year), yet for all other multiples of four. This figure shows a histogram of the strongest seasonalities in each feature.
@@ A.3 Details on Model Configurations and Training
We based our analysis on the implementation in the darts software library (Herzen et al., 2022) and the official implementation of Chronos (Ansari et al., 2024) [^1] https://github.com/amazon-science/chronos-forecasting. We make our implementation openly available [^2] https://github.com/felixdivo/Forecasting-Company-Fundamentals. We used scikit-learn (Pedregosa et al., 2011) for data normalization. For evaluation, we used torchmetrics (Nicki Skafte Detlefsen et al., 2022) and the CRPS implementation of Pyro (Bingham et al., 2019). Some models obtain probabilistic forecasts via specialized architectures or sampling methods, while others simply learn to predict the parameters of a distribution instead of the data directly. This is the case for Linear Regression as well as all deep learning models, where we, therefore, performed quantile regression on $0.01,0.05,0.1,0.15,0.2,0.25,0.3,0.4,0.5,0.6,0.7,0.75,0.8,0.85,0.9,0.95,0.99$.
@@
For the linear regression and random forest models, we estimated separate models for each target variable and future time step. For the latter, each forest contained 100 trees learned from bootstrapped subsamples.
We trained all deep learning models with gradient descent on a batch size of 64 for 100 epochs. We adjusted for different training duration requirements by early stopping after no improvement in the validation loss after three epochs. We employed the AdamW optimizer (Loshchilov & Hutter, 2018) with learning rate $10^{-4}$, weight decay weighting of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. Furthermore, we clipped gradients to $1.0$ to stabilize training. For all models except DLinear and NLinear, we used a 10% dropout rate. The following lists the specific hyperparameters for each of the deep learning architectures. For DLinear, we used a kernel size of 10 to estimate the moving average of the trend. All recurrent neural networks (GRU, LSTM, and block variants) were three layers deep. The single models and block variants had a hidden size of 64 and 128, respectively. The TCN convolutions were set to a kernel width of 3 steps, 16 filters, and a dilation of 2 time steps. The Transformer was trained with a token size of 120, a feedforward dimension of 512, GELU activations, four encoder layers, four decoder layers, and six attention heads. TFT used a hidden size of 36 and six attention heads covering the entire time span. We learned a single block of six layers with hidden dimension 512 for N-BEATS. The N-HiTS models consisted of three stacks, each with a single two-layer block and 512 hidden dimensions. Our configuration of TiDE had two encoder and decoder layers each, a hidden size of 128, and a decoder with a hidden size of 32 and an output dimension of 16.
A.4 On the Use of Reversible Instance Norm
@@
Table 5: RevIN is beneficial for only some of the models. The scores are the nCRPS obtained from probabilistic forecasts, where lower is better.
@@
A.5 On the Use of Static Features
As shown in Table 1, some models can leverage numerical or categorical context in addition to the plain time series. As laid out in Section 4.1.1, we added GICS sector classifications to each company. In this section, we investigate the impact of this additional information on forecast accuracy. As Table 6 shows, the overall effect of this additional data is marginal. While it slightly improved the performance of NLinear, it worsened the performance of TFT, conceivably due to the larger tendency to overfit.
Table 6: Most models perform similarly with and without incorporating static features. This table compares supplying the models with only the time series to also providing static context per company. The scores are the nCRPS obtained from probabilistic forecasts, where lower is better.
Models Without Statics With Statics Improvement
Linear Reg. 0.401±0.02 0.401±0.02 0.00%
DLinear 0.131±0.02 0.131±0.02 0.00%
NLinear 0.136±0.02 0.135±0.02 0.74%
TFT 0.132±0.02 0.134±0.02 -1.52%
TiDE 0.126±0.02 0.126±0.02 0.00%
A.6 Algorithmic and Human Forecasts
To facilitate a relation of the human evaluation to the automated forecast, we provide comparable results in Table 7, as discussed in Section 4.1.3. Here, we made sure to test the models only on the companies where expert forecasts were available. Furthermore, only the fourth (one-year-ahead) time step is considered, which aligns with the human expectations. For simplicity, we show the results for an entire company in the model evaluation, even if individual human analyst predictions are missing, since this happens only rarely. Since the expectations contained some heavy outliers, we omitted metric results larger than four standard deviations from the statistics. The comparison is performed on the transformed data as laid out in Section 4.1.2. Consistent with the prior evaluation, the provided nCRPS score for the deterministic analyst expectations is the special case of the MAE. Results are shown for the probabilistic models.
@@ A.6 Algorithmic and Human Forecasts
xLSTM-Mixer 0.200±0.03 0.383±0.14 0.225±0.03 3.045±1.27 7528.145 0.422±0.02 -27315.441 0.141±0.02
Chronos Small 0.217±0.03 0.649±0.40 0.255±0.04 3.442±1.51 7968.601 0.421±0.02 -27482.442 0.409±0.02
Chronos Base 0.224±0.03 0.686±0.34 0.264±0.04 3.425±1.28 8627.039 0.423±0.02 -27558.638 0.409±0.02